# Dual-Polarized Intelligent Omni-Surfaces for Independent Reflective-Refractive Transmission

### v1

@@ IV-A Analysis of the XPD
Given the optimal power allocation based on (23) and (24), we first analyze the impact of polarization leakage, i.e., XPD on the digital beamformer ${\bf W}=[{\bf W_{r}},{\bf W}_{t}]$.
Proposition 1: If the polarization leakage is zero, each MU’s digital beamformer ${\bf W}^{(k)}$ will have half of its elements corresponding to encoding elements in the opposite polarization direction, with values equal to zero.
@@
Proof: Note that when ${\beta}_{\rm BI}=0$, the polarization of the transmitted signal will not change when it propagates from the BS. Alternatively, ${\beta}_{\rm BI}=1$ correspond to the case that the transmitted signal will be completely converted to the opposite polarization and can also be decomposed as two dependent subchannels. On the other hand, if $10<\beta_{\rm BI}<$ 1, a part of the power will be leaked to the other polarization, which cause cross-polarization interference and degrade the performance.
Based on Proposition 2, we can obtain Proposition 3.
Proposition 3: Low polarization leakage is not always better for the dual-polarized IOS-assisted systems. High leakage is beneficial for a low transmit power.
Proof: The system performance initially decreases and then increases as $\beta_{\rm BI}$ increases. In some cases, the worst performance may occur at a smaller value of $\beta_{\rm BI}$, which implies that a larger $\beta_{\rm BI}$ could potentially help improve performance.
IV-B Analysis of the user ratio ${\mathcal{N}}_{r}/{\mathcal{N}}_{t}$
In this section, we analyze the impact of the user ratio of the reflection and refraction on the sum-rate of an IOS-assisted communication system.
Proposition 4: Different from traditional IOS, the sum rate of dual-polarized IOS-assisted systems are not impacted by the user ratio ${\mathcal{N}}_{r}/{\mathcal{N}}_{t}$.
Proof: The sum rate of traditional IOS-assisted systems is impacted by ${\mathcal{N}}_{r}/{\mathcal{N}}_{t}$, since the elements have a fixed power ratio. When the power ratio is unmatched with the user ratio on both sides, the performance is poor. A dual-polarized IOS-assisted system can overcome this by power allocation for different polarized antennas at the BS. Thus, the sum rate is not impacted by the ratio ${\mathcal{N}}_{r}/{\mathcal{N}}_{t}$.
@@ V Simulation Results
In this section, we evaluate the performance of the considered dual-polarized IOS-assisted system based on the proposed algorithm and compare it with an RIS-aided system, a traditional IOS-aided system, and a conventional cellular system in the absence of IRS or IOS. In the conventional cellular system, the MUs receive only the direct links from the BS without the assistance of a reconfigurable surface.
Figure 2: Sum rate versus XPD with different transmit power.
In Figure. 2, we plot the sum rate versus XPD for different transmit power. It can be observed that as the XPD increases from 0 to 1, the communication rate for different transmit powers initially increases and then decreases, which is consistent with our previous analysis. Furthermore, it is noticeable that the location of the peak is related to the transmit power.
Figure 3: Sum rate versus transmit power with different XPD.
Figure 4: Sum rate versus transmit power with different user ratio.
In Figure. 3, we plot the sum rate versus the transmit power for different XPD factors. It can be observed that as the transmit power increases, both the system and communication rate improve for the two curves. However, when the XPD factor is 0.1, the communication rate is lower at lower transmit powers, whereas at higher transmit powers, the communication rate is noticeably higher than when the XPD factor is 0.2. This further validates the previous analysis that a lower XPD is not always beneficial.
In Figure. 4, we plot the sum rate versus the ratio of the reflection user and refraction user for different scenarios. We have demonstrated the impact of the user ratio on system performance in MISO communication systems with dual-polarized IOS, traditional IOS, and without any IOS assistance. For the traditional IOS, since the transmission-to-reflection ratio is fixed, the system performs well when the user ratio on both sides aligns well with the transmission-reflection ratio. However, in extreme cases, the performance suffers significantly. In contrast, for the polarization IOS, it can allocate signal power to the two sides of the metasurface via precoding at the base station. As a result, the performance is better compared to the MISO system without IOS assistance, and the system performance is less sensitive to the user ratio.
@@ VI Conclusion
In this letter, we considered a dual-polarized IOS-aided MISO downlink system and investigated the joint BS digital and ios analog beamforming design to maximize the sum rate for independent reflective-refractive transmission by utilizing the polarization domain. Then, we analyzed the impact of the XPD and the user ratio on both sides of the dual-polarized IOS on the system performance. We have the following conclusions: 1)Unlike traditional IOS, the performance of the polarization IOS is less affected by the ratio of the number of users on both sides, as the coupling between transmission and reflection coefficients is decoupled. 2)Low polarization is not always better for the dual-polarized IOS-aided system.

### v2

@@ IV-A Analysis of the XPD
The power allocated to the vertical and horizontal polarization is expressed as
$p_{v}={\rm Tr}(\frac{\sum_{k=1}^{K_{r}}|u_{r,k_{r}}f_{r,k_{r}}|^{2}{\bf h}_{r,k_{r}}^{\rm H}{\bf h}_{r,k_{r}}}{({\bf M}+\lambda{\bf I})^{2}}),$ (28)
$p_{h}={\rm Tr}(\frac{\sum_{k=1}^{K_{t}}|u_{t,k_{t}}f_{t,k_{t}}|^{2}{\bf h}_{t,k_{t}}^{\rm H}{\bf h}_{t,k_{t}}}{({\bf M}+\lambda{\bf I})^{2}}).$ (29)
It can be noted that the factor $\lambda$ does not affect the power distribution between the two polarization directions. Polarization leakage caused by the channel is what affects the power distribution. Then, we analyze the impact of polarization leakage, i.e., XPD on the digital beamformer ${\bf W}=[{\bf W}_{r},{\bf W}_{t}]$.
Proposition 1: If the polarization leakage is zero, each MU’s digital beamformer ${\bf W}^{(k)}$ will have half of its elements corresponding to the antennas in the opposite polarization direction equal to zero.
@@
Proof: Note that when ${\beta}_{\rm BI}=0$, the polarization of the transmitted signal will not change when it propagates from the BS. Alternatively, ${\beta}_{\rm BI}=1$ corresponds to the case that the transmitted signal will be completely converted to the opposite polarization and can also be decomposed as two dependent subchannels. On the other hand, if $10<\beta_{\rm BI}<$ 1, a part of the power will be leaked to the other polarization, which causes cross-polarization interference and degrades the performance.
Based on Proposition 2, we can obtain Remark 1.
Remark 1: High polarization leakage may provide extra gains for dual-polarized IOS-assisted systems.
IV-B Analysis of the user ratio ${K}_{r}/{K}_{t}$
The sum rate of power-domain IOS-assisted systems is impacted by ${K}_{r}/{K}_{t}$ since the elements have a fixed power ratio. When the power ratio is unmatched with the user ratio on both sides, the performance is poor. A dual-polarized IOS-assisted system can overcome this by power allocation for different polarized antennas at the BS.
Remark 2: The sum rate of dual-polarized IOS-assisted systems is not impacted by the user ratio ${K}_{r}/{K}_{t}$.
Figure 2: Sum rate versus $K_{r}/(K_{t}+K_{r})$.
Figure 3: Sum rate versus XPD.
Figure 4: Sum rate versus transmit power.
Proposition 4: When the number of users on both sides of the IOS is the same, symmetrically distributed, and the power allocation ratio is 1, the direct link leads to a performance gap between the power-domain IOS and the dual-polarized IOS.
Proof: Considering a power-domain IOS with power-split ratio $\epsilon$, the sum rate of the MUs on both sides is expressed as
$\sum_{k=1}^{K_{r}+K_{t}}R_{k}=\sum_{{k_{r}\in\mathcal{N}}_{r}}$ ${\log_{2}(1+\tau_{k_{r}}+\frac{\epsilon}{1+\epsilon}\chi_{k_{r}})}$
$+\sum_{k_{t}\in{\mathcal{N}}_{t}}{\log_{2}(1+\tau_{k_{t}}+\frac{1}{1+\epsilon}\chi_{k_{t}})},$ (30)
where $\tau_{k}$ is the SINR of ${k}^{th}$ MU from direct link, $\chi_{k}$ is the SINR of $k^{th}$ MU from the IOS-aided link. By taking the first derivative of (IV-B) with respect to $\epsilon$ , we obtain
$\sum_{k=1}^{K_{r}+K_{t}}\frac{dR_{k}}{d\epsilon}=\frac{1}{\rm ln2}$ $\sum_{k_{t}\in{\mathcal{N}}_{t}}{\frac{\chi_{k_{t}}/(1+\epsilon)^{2}}{1+\tau_{k_{t}}+\frac{\epsilon}{1+\epsilon}\chi_{k_{t}}}}$
$-\frac{1}{\rm ln2}\sum_{k_{r}\in{\mathcal{N}}_{r}}{\frac{\chi_{k_{r}}/(1+\epsilon)^{2}}{1+\tau_{k_{r}}+\frac{1}{1+\epsilon}\chi_{k_{r}}}}.$ (31)
Here, we assume $\chi_{k_{t}}=\chi_{k_{r}}=\chi_{k}$. If $\tau_{k_{r}}=\tau_{k_{t}}$, we have $\epsilon^{opt}=1$ and there is no performance gap between the power-domain and dual-polarized IOS. Otherwise, $\epsilon^{opt}\neq 1$ and a performance gap exists.
@@ V Simulation Results
In this section, we evaluate the performance of the considered dual-polarized IOS-assisted system based on the proposed algorithm and compare it with a power-domain IOS-aided system, a dual-polarized RIS-aided system and a cellular system. In the conventional cellular system, the MUs receive only the direct links from the BS without the assistance of an IOS. The details of the simulation parameters are as follows: number of BS antennas $N_{t}=4$, number of IOS elements $M=4$, number of MUs $K_{r}=K_{t}=2$, XPD factors $\beta_{\rm BI}=\beta_{\rm IU}=\beta_{\rm BU}=0.1$, the transmit power $P_{\rm BS}=1{\rm W}$. The values of XPD are chosen according to [10].
In Figure. 4, we plot the sum rate versus ${K}_{r}/({K}_{r}+K_{t})$. For the power-domain IOS, it performs well when $K_{r}/K_{t}$ aligns well with the fixed $\epsilon$. In this case, we assume the MUs are symmetrically located and $\tau_{k_{r}}=\tau_{k_{t}}$. According to Proposition 4, there is no performance gap between the dual-polarized IOS and power-domain IOS when $K_{r}=K_{t}$. However, in some extreme cases, such as when the x-axis coordinate is 0.2 or 0.8, power is wasted due to the mismatch between $\epsilon$ and the number of MUs, resulting in a performance gap of 1.2 bit/s/Hz. In addition, the dual-polarized RIS-aided system performs well as ${K}_{r}/({K}_{r}+K_{t})$ gradually increases since it can only reflect the signal.
In Figure. 4, we present the sum rate versus different XPD factors, considering various numbers of BS antennas and IOS elements. The results show that as the XPD factor between the BS and the IOS, $\beta_{\rm BI}$, increases from 0 to 1, the sum rate first decreases and then increases, which aligns with Proposition 2. Furthermore, increasing the number of IOS elements and transmit antennas, compared to the benchmark configuration of $N_{t}=2,M=4,K=2$, results in improved performance.
In Figure 4, the sum rate is plotted versus transmit power for varying XPD factors. As transmit power increases, the sum rate improves due to a higher SINR. Increasing the $\beta_{\rm BI}$ from 0.1 to 0.2 results in performance degradation. This is due to increased polarization leakage, which not only reduces signal power but also increases cross-polarization interference, lowering the SINR. When $\beta_{\rm BI}$ increases to 0.8, the performance surpasses that of the other cases. This is because the zero elements described in Proposition 1 increase, leading to more power being allocated to the orthogonal polarization. These results are consistent with the analysis in Remark 1.
@@ VI Conclusion
In this paper, we considered a dual-polarized IOS-aided downlink system and investigated the joint BS digital and IOS analog beamforming design to maximize the sum rate for reflective-refractive transmission by utilizing the polarization domain. Then, we analyzed the impact of the XPD and the user ratio on both sides of the dual-polarized IOS. We have the following conclusions: 1) Unlike power-domain IOS, the performance of the dual-polarized IOS is less affected by the ratio of the number of MUs on both sides, as the coupling between refraction and reflection coefficients is decoupled. 2) High polarization leakage may provide extra gains for dual-polarized IOS-assisted systems.